OpenStudy (savy):

Write an equation of a line that passes through (-9,3) and is perpendicular to y+3x=-17

OpenStudy (mayankdevnani):

y+3x=-17 can you identify the slope of this given line ?

OpenStudy (mayankdevnani):

any ideas ?

OpenStudy (savy):

the slope of that one is -3 so for the perpendicular line it has to be 1/3

OpenStudy (mayankdevnani):

good!

OpenStudy (mayankdevnani):

so equation of perpendicular line would be :- \[\large \bf y-y_1=\frac{1}{3}(x-x_1)\] where,\(\large \bf (x_1,y_1)\) is point through which perpendicular line would pass

OpenStudy (mayankdevnani):

can you tell me what is (x1,y1) ?

OpenStudy (adjunctbear):

Look at the points (-9,3) Consider which is x and which is y and voila you will have X1 and y1!